# arXiv:2512.09295v2[math.ST] 01 Mar 2026

### v2

@@ Abstract
We revisit the signal denoising problem through the lens of optimal transport: the goal is to recover an unknown scalar signal distribution $X\sim P$ from noisy observations $Y=X+\sigma Z$, with $Z$ being standard Gaussian independent of $X$ and $\sigma>0$ a known noise level. Let $Q$ denote the distribution of $Y$. We introduce a hierarchy of denoisers $T_{0},T_{1},\ldots,T_{\infty}:\mathbb{R}\to\mathbb{R}$ that are agnostic to the signal distribution $P$, depending only on higher-order score functions of $Q$. Each denoiser $T_{K}$ is progressively refined using the (2K1)(2K-1)-th order score function of $Q$ at noise resolution $\sigma^{2K}$, achieving better denoising quality measured by the Wasserstein metric $W(T_{K}\sharp Q,P)$. The limiting denoiser $T_{\infty}$ identifies the optimal transport map with $T_{\infty}\sharp Q=P$.
We provide a complete characterization of the combinatorial structure underlying this hierarchy through Bell polynomial recursions, revealing how higher-order score functions encode the optimal transport map for signal denoising. We study two estimation strategies with convergence rates for higher-order scores from i.i.d. samples drawn from $Q$: (i) plug-in estimation via Gaussian kernel smoothing, and (ii) direct estimation via higher-order score matching. This hierarchy of agnostic denoisers presents new opportunities for signal denoising and empirical Bayes research.
Keywords— Agnostic denoisers, higher-order score functions, distributional shrinkage, optimal transport, Bell polynomials, advanced combinatorics.
@@ 1 Introduction
Let $X\in\mathbb{R}$ be an unobserved signal drawn from an unknown distribution $P$, and $Z$ be a standard Gaussian noise independent of $X$. We only observe the noisy measurements $Y$ from an additive model, with noise level $\sigma\in\mathbb{R}_{>0}$ known a priori:
@@
We denote the distribution of $Y$ as $Q$, and denote its density function with respect to the Lebesgue measure as $q:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$.
@@ 1.1 Stein’s Paradox and Shrinkage
Stein’s paradox [JS61, Bro66, EM73] states the fact that it is possible to estimate $X$ from $Y$ with a smaller mean squared error (MSE) than $Y$ itself. Consider the estimator $Y+h(Y)$, where $h:\mathbb{R}\rightarrow\mathbb{R}$ is a denoising function. Stein’s unbiased risk estimate (SURE) [Ste81] gives an unbiased estimate of the MSE:
@@
[…]
@@
[…]
$=\sigma^{2}-\sigma^{4}\mathop{\mathbb{E}}\left(\frac{q^{\prime}(Y)}{q(Y)}\right)^{2}<\sigma^{2}$ $\text{when}~h(y)=\sigma^{2}\frac{q^{\prime}(y)}{q(y)}$
@@
The other notable example is the James-Stein estimator. Take $P=\frac{1}{d}\sum_{j=1}^{d}\delta_{X_{j}}$ as a discrete distribution with $d\geq 2$ atoms. For $j=1,\ldots,d$, let $Z_{j}$ be independent standard Gaussian noise and $Y_{j}=X_{j}+\sigma Z_{j}$. With the vectors noted by $\mathbf{Z}=(Z_{1},\ldots,Z_{d})$ and $\mathbf{Y}=(Y_{1},\ldots,Y_{d})$, the James-Stein estimator [JS61] is equivalent to the component-wise denoiser,
@@
$\textrm{MSE}(h):=\frac{1}{d}\sum_{j=1}^{d}\mathop{\mathbb{E}}_{\mathbf{Z}}[|Y_{j}+h_{\mathbf{Y}_{-j}}(Y_{j})-X_{j}|^{2}]$
$=\sigma^{2}+\frac{1}{d}\sum_{j=1}^{d}\mathop{\mathbb{E}}_{\mathbf{Z}}\left[h_{\mathbf{Y}_{-j}}^{2}(Y_{j})+2\sigma^{2}\partial_{Y_{j}}h_{\mathbf{Y}_{-j}}(Y_{j})\right]$ same as Equation (1.1)
$=\sigma^{2}+\sigma^{4}\mathop{\mathbb{E}}_{\mathbf{Z}}\left[\frac{(d-2)^{2}}{d}\sum_{j=1}^{d}\frac{Y_{j}^{2}}{(\|\mathbf{Y}\|^{2})^{2}}-2\frac{(d-2)}{d}\sum_{j=1}^{d}\left(\frac{1}{\|\mathbf{Y}\|^{2}}-\frac{2Y_{j}^{2}}{(\|\mathbf{Y}\|^{2})^{2}}\right)\right]$
$=\sigma^{2}-\sigma^{4}\frac{(d-2)^{2}}{d}\mathop{\mathbb{E}}_{\mathbf{Z}}\left[\frac{1}{\|\mathbf{Y}\|^{2}}\right]<\sigma^{2}$ $\text{when}~d\geq 3$
@@
Why the Wasserstein metric? In the case when $X\sim\frac{1}{d}\sum_{j=1}^{d}\delta_{X_{j}}$, a discrete distribution,
$W_{2}^{2}(\mathcal{L}(Y+h(Y)),\mathcal{L}(X))=\min_{\varsigma\in S_{d}}\frac{1}{d}\sum_{j=1}^{d}\mathop{\mathbb{E}}[|Y_{j}+h(Y_{j})-X_{\varsigma(j)}|^{2}]\leq\frac{1}{d}\sum_{j=1}^{d}\mathop{\mathbb{E}}[|Y_{j}+h(Y_{j})-X_{j}|^{2}]$
where $\varsigma\in S_{d}$ is a permutation. More generally, for any random variables $X,Y$,
@@
where the inequality follows by considering better couplings beyond the naive coupling $(Y+h(Y),X)$. Allowing for permutations, or more generally, couplings, offers leeway for higher denoising accuracy, but, perhaps more importantly, it also opens up new theoretical and practical perspectives. In theory, it uncovers new mathematical structures from advanced combinatorics [Com74], optimal transport [Vil08], and higher-order score functions [Bob24]. In practice, for several modern applications in generative modeling, such as image denoising via score-based diffusion models [Son+20, Soh+15, LDK26], the Wasserstein metric is more relevant than the MSE; people care about the quality of the denoiser at the distributional level, not just at the data level.
Traditional data-level shrinkage methods including Bayes-optimal and empirical Bayes denoisers [EM73, JS61] often over-shrink the distribution. They produce denoised distributions that are overly concentrated and do not match the actual signal distribution [GS24, JIS25, Lia25]. We refer readers to [Lia25, JIS25] for examples and discussions on the benefits of distribution-level shrinkage compared to traditional data-level shrinkage. This paper continues along the path of [GS24, JIS25, Lia25] in studying classic signal denoising through the lens of optimal transport.
@@ 1.2 A Hierarchy of Optimal Transport Denoisers
In this paper, we study signal denoising under the Wasserstein metric and examine whether shrinkage helps denoising in a distributional sense. Recall that for $r\geq 1$, the $r$-Wasserstein metric [Vil03, Theorem 2.18] on the real line takes the closed-form expression,
[…]
@@
Consider the denoiser of the form $Y+h(Y)$, we look for denoising function $h:\mathbb{R}\rightarrow\mathbb{R}$ that achieves significantly smaller error than the trivial denoiser $Y$, measured by the Wasserstein metric, namely,
@@
This is analogous to Stein’s paradox, but now at the distributional level. Under mild regularity conditions, the optimal denoiser in the Wasserstein sense is the unique optimal transport map pushing $Q$ onto $P$ [Vil03], denoted as $T_{\infty}:\mathbb{R}\to\mathbb{R}$
@@
To see how such a hierarchy is constructed, we note that for a monotonically increasing map $y\mapsto y+h(y)$
[…]
@@
These formulas may look mysterious, but they follow a combinatorial relationship involving Bell polynomials; see Theorem 3 for the general formula for $h_{k},k\geq 1$:
$h_{k}=-(-1)^{k}\frac{G^{(2k)}}{G^{(1)}}-\sum_{i=1}^{k-1}(-1)^{k-i}\binom{k}{i}\sum_{j=0}^{i-1}\frac{G^{(2k-i-j)}}{G^{(1)}}B_{i,i-j}(h_{1},\ldots,h_{j+1})-\sum_{j=0}^{k-2}\frac{G^{(k-j)}}{G^{(1)}}B_{k,k-j}(h_{1},\ldots,h_{j+1})\;.$
Here $B_{n,k}$’s are the partial Bell polynomials [Com74] as in Definition 3, capturing the combinatorial structure of integer partitions. The above recursions can be solved sequentially to obtain $h_{k},k\geq 1$.
Curiously, we emphasize that such denoising functions $h_{k},k\geq 1$ only depend on higher-order score functions of $Q$, and are agnostic to the specific form or knowledge of the signal distribution $P$. This phenomenon motivates the name agnostic denoisers. This is in the spirit of the James-Stein denoiser, which is also agnostic to the prior distribution $P$; however, the James-Stein denoiser is only agnostic for discrete distributions with $d$ atoms. In contrast, our denoisers $T_{K},K\geq 1$ are agnostic among all distributions on $\mathbb{R}$.
@@
The paper is organized as follows. As a warm-up for the technical combinatorial structure, Section 2.1 and Theorem 1 study the infinite expansion of the optimal transport map $T_{\infty}=F^{-1}\circ G$ in terms of higher-order score functions of the signal distribution $P$. Section 2.2 and Theorem 2 quantify $W_{r}(T_{K}\sharp Q,P)$, the denoising accuracy of $T_{K},K\geq 1$, as well as the uniform error of the map $T_{K}$ in approximating the optimal transport map $T_{\infty}=F^{-1}\circ G$. Section 2 is in reminiscence of g-modeling in empirical Bayes [Efr14], where one uses prior information to construct denoisers. Though the results in Section 2 are primarily of theoretical interest, they prepare the technical tools for the main results in Section 3.
Moving on to the main results, Section 3 derives and studies optimal transport denoisers via higher-order score functions of $Q$, directly identifiable and estimable based on observations $Y$. Section 3.1 and Theorem 3 represent $\{h_{k}\}_{k=1}^{\infty}$ as polynomials of higher-order score functions $\{\tfrac{q^{(m)}}{q}\}_{m=1}^{\infty}$. This result reveals how higher-order score functions encode information of the optimal transport denoising map, a finding new to the literature. We then study estimation strategies and rates of convergence for these higher-order scores, based on $n$ i.i.d. samples $\{Y_{i}\}_{i=1}^{n}$. We present two approaches with theoretical guarantees on the estimation accuracy: (i) a plug-in estimation approach via Gaussian kernel smoothing that estimates $q^{(m)}(y)$ locally at any given $y$, see Section 3.2 and Theorem 4 for the rate of convergence; (ii) a direct estimation approach via higher-order score matching that estimates the function $\tfrac{q^{(m)}}{q}$ globally, see Section 3.3 and Theorem 5 for the rate of convergence. This is in reminiscence of f-modeling in empirical Bayes [Efr14].
@@ 1.3 Relations to the Literature
The optimal transport (OT) denoisers are fundamentally different from the Bayes-optimal or empirical Bayes denoisers [EM73, Rob64, JS61] that minimize the mean squared error. To wit, the Bayes-optimal denoiser takes the form
@@
different from the hierachy of OT denoisers $T_{K},K\geq 0$: with $T_{0}(y)=y$, and
@@
Denoiser $T_{1}$ has been studied extensively in diffusion models; see [LDK26, BB25]. [Lia25] studies these two OT denoisers $T_{1},T_{2}$ in a general $d$-dimensional setting, and establishes that these two denoisers are universal denoisers. They remain universal to a wide range of noise distributions, extending beyond the Gaussian setting that requires Tweedie’s formula. On the one hand, this current paper generalizes the results in [Lia25] and derives a complete hierarchy of OT denoisers to arbitrary higher-order $K$. On the other hand, this paper considers only $d=1$ and the Gaussian noise.
@@
Empirical Bayes methods [EM73, Rob64, JS61, Gho+25] study the Bayes-optimal denoiser constructed based on i.i.d. samples $\{Y_{i}\}_{i=1}^{n}$ drawn from $Q$. There are two distinct approaches, named the f-modeling (modeling on the observed data, the $Y$ space), and g-modeling (modeling on the unobserved prior/signal, or the $X$ space) [Efr14]; we refer readers to [IS25, Chapter 6] for detailed discussions. [Efr14] states that g-modeling on the $X$ space has dominated the theoretical literature in empirical Bayes, see [Efr16, JZ09, Lai78] for examples. In particular, [JZ09] studied nonparametric maximum likelihood estimation (NPMLE) of the prior distribution $P$ based on i.i.d. samples $\{Y_{i}\}_{i=1}^{n}$ drawn from $Q$. They established the rate of convergence of NPMLE $\widehat{P}$ to the true prior $P$ in terms of the Hellinger distance between the induced densities on the $Y$ space. [Gho+25] draws the connection between score matching [HD05] and SURE [Ste81, XKB12] to propose a SURE training procedure; they also establish the corresponding rate of convergence to the true prior measured by Fisher divergence. We note that the theoretical empirical Bayes literature primarily focuses on estimating the prior $P$ first, then plugging in Tweedie’s formula to construct the empirical Bayes denoiser. However, as noted in [Efr14], the f-modeling directly on the $Y$ space, is more natural for constructing denoisers, since the data $\{Y_{i}\}_{i=1}^{n}$ are directly observed. [Efr14] proposed several parametric models on the $Y$ space to estimate the Bayes-optimal denoiser.
@@
This paper approaches this denoising question in a drastically different way from the empirical Bayes literature: we construct denoisers directly on the observational $Y$ space via higher-order score functions, without estimating the prior $P$ on the $X$ space. The denoisers $T:\mathbb{R}\rightarrow\mathbb{R}$ constructed in this paper are optimal in the Wasserstein sense, and when applied to $Y\sim Q$, the denoised distribution $T\sharp Q$ matches the signal distribution $P$ up to high accuracy. In other words, this paper proposes directly estimating the optimal transport map $T$ that pushes $Q$ onto $P$, without the knowledge or estimation of $P$. This is in sharp contrast to the g-modeling approach, which first estimates $P$, then uses that information to construct empirical Bayes denoisers. Our denoisers are nonparametric in nature, agnostic to the specific form of the prior distribution $P$. In terms of estimation of the higher-order score functions, we present two approaches: (i) a plug-in estimation via Gaussian kernel smoothing; (ii) a direct estimation approach via higher-order score matching. Both approaches come with theoretical guarantees on the rates of convergence. Crucially, our optimal transport approach discovers and isolates a notion of agnostic denoisers: the denoisers $T_{K},K\geq 1$ are agnostic to the specific form of $P$; they only depend on the higher-order score functions of $Q$.
@@
Score matching techniques have received considerable attention in the machine learning literature [HD05, Hyv07, Vin11, SH19]. The traditional score matching focuses on estimating the first-order score function $\tfrac{q^{(1)}}{q}$ [Hyv07, Sar+18, Sri+17, Fen+24] via Fisher divergence. [Bob24] recently studied Fisher-type information induced by higher-order score functions $\tfrac{q^{(m)}}{q},m\geq 1$. We establish estimation strategies and rates of convergence for these higher-order scores, and derive through advanced combinatorics that these higher-order scores encode full information to construct the OT denoising map with $T\sharp Q=P$. To our knowledge, this paper is the first to fully characterize the optimal transport denoisers in terms of higher-order score functions; we make rigorous the connection among information geometry, optimal transport, and advanced combinatorics in this classic denoising problem.

### v3

# Distributional Shrinkage II: Higher-Order Scores

@@ Abstract
Consider the additive Gaussian model $Y=X+\sigma Z$, where $X\sim P$ is an unknown signal, $Z\sim\mathcal{N}(0,1)$ is independent of $X$, and $\sigma>0$ is known. Let $Q$ denote the law of $Y$. We construct a hierarchy of denoisers $T_{0},T_{1},\ldots,T_{\infty}\colon\mathbb{R}\to\mathbb{R}$ that depend only on higher-order score functions $q^{(m)}/q$, $m\geq 1$, of $Q$ and require no knowledge of the law $P$. The $K$-th order denoiser $T_{K}$ involves scores up to order 2K12K{-}1 and satisfies $W_{r}(T_{K}\sharp Q,P)=O(\sigma^{2(K+1)})$ for every $r\geq 1$; in the limit, $T_{\infty}$ recovers the monotone optimal transport map (Brenier map) pushing $Q$ onto $P$.
We provide a complete characterization of the combinatorial structure governing this hierarchy through partial Bell polynomial recursions, making precise how higher-order score functions encode the Brenier map. We further establish rates of convergence for estimating these scores from $n$ i.i.d. draws from $Q$ under two complementary strategies: (i) plug-in kernel density estimation, and (ii) higher-order score matching. The construction reveals a precise interplay among higher-order Fisher-type information, optimal transport, and the combinatorics of integer partitions.
Keywords— optimal transport denoisers, higher-order score functions, distributional shrinkage, Gaussian deconvolution, Monge-Ampère equation, Bell polynomials, Brenier map.
@@ 1 Introduction
Let $X\in\mathbb{R}$ be an unobserved signal drawn from an unknown distribution $P$, and let $Z$ be a standard Gaussian random variable independent of $X$. We observe only the noisy measurement $Y$ from an additive model, with noise level $\sigma\in\mathbb{R}_{>0}$ known a priori:
@@
We denote the distribution of $Y$ by $Q$ and its density function with respect to the Lebesgue measure by $q:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$.
@@ 1.1 Stein’s Paradox and Shrinkage
Stein’s paradox [JS61, Bro66, EM73] states that it is possible to estimate $X$ from $Y$ with a smaller mean squared error (MSE) than $Y$ itself. Consider the estimator $Y+h(Y)$, where $h:\mathbb{R}\rightarrow\mathbb{R}$ is a denoising function. Stein’s unbiased risk estimate (SURE) [Ste81] gives an unbiased estimate of the MSE:
@@
$=\sigma^{2}+\mathop{\mathbb{E}}[h^{2}(Y)]+2\sigma^{2}\mathop{\mathbb{E}}[h^{\prime}(X+\sigma Z)]$
$\hskip 18.49988pt\left(\text{by integration by parts:}~\textstyle\int ze^{-z^{2}/2}h(x+\sigma z)\,\mathrm{d}z=\sigma\int h^{\prime}(x+\sigma z)e^{-z^{2}/2}\,\mathrm{d}z\right)$
@@
$=\sigma^{2}+\int q(y)\left(h^{2}(y)-2\sigma^{2}\frac{q^{\prime}(y)}{q(y)}h(y)\right)\,\mathrm{d}y$
$\hskip 18.49988pt\left(\text{by integration by parts:}~\textstyle\int h^{\prime}(y)q(y)\,\mathrm{d}y=-\int h(y)q^{\prime}(y)\,\mathrm{d}y\right)$
$=\sigma^{2}-\sigma^{4}\mathop{\mathbb{E}}\left(\frac{q^{\prime}(Y)}{q(Y)}\right)^{2}<\sigma^{2}\;,\hskip 18.49988pt\text{when}~h(y)=\sigma^{2}\frac{q^{\prime}(y)}{q(y)}\;.$
@@
The other notable example is the James-Stein estimator. Take $P=\frac{1}{d}\sum_{j=1}^{d}\delta_{X_{j}}$ as a discrete distribution with $d\geq 2$ atoms. For $j=1,\ldots,d$, let $Z_{j}$ be independent standard Gaussian noise and $Y_{j}=X_{j}+\sigma Z_{j}$. With the vectors denoted by $\mathbf{Z}=(Z_{1},\ldots,Z_{d})$ and $\mathbf{Y}=(Y_{1},\ldots,Y_{d})$, the James-Stein estimator [JS61] is equivalent to the component-wise denoiser,
@@
$\textrm{MSE}(h)$ $:=\frac{1}{d}\sum_{j=1}^{d}\mathop{\mathbb{E}}_{\mathbf{Z}}\!\left[|Y_{j}+h_{\mathbf{Y}_{-j}}(Y_{j})-X_{j}|^{2}\right]$
[…]
$=\sigma^{2}+\sigma^{4}\mathop{\mathbb{E}}_{\mathbf{Z}}\Bigg[\frac{(d-2)^{2}}{d}\sum_{j=1}^{d}\frac{Y_{j}^{2}}{\|\mathbf{Y}\|^{4}}-\frac{2(d-2)}{d}\sum_{j=1}^{d}\left(\frac{1}{\|\mathbf{Y}\|^{2}}-\frac{2Y_{j}^{2}}{\|\mathbf{Y}\|^{4}}\right)\Bigg]$
$=\sigma^{2}-\sigma^{4}\frac{(d-2)^{2}}{d}\mathop{\mathbb{E}}_{\mathbf{Z}}\left[\frac{1}{\|\mathbf{Y}\|^{2}}\right]<\sigma^{2}\;,\hskip 18.49988pt\text{when}~d\geq 3\;.$
@@
Why the Wasserstein metric? In the case where $X\sim\frac{1}{d}\sum_{j=1}^{d}\delta_{X_{j}}$, a discrete distribution,
$W_{2}^{2}\big(\mathcal{L}(Y+h(Y)),\mathcal{L}(X)\big)=\min_{\varsigma\in S_{d}}\frac{1}{d}\sum_{j=1}^{d}\mathop{\mathbb{E}}\left[|Y_{j}+h(Y_{j})-X_{\varsigma(j)}|^{2}\right]\leq\frac{1}{d}\sum_{j=1}^{d}\mathop{\mathbb{E}}\left[|Y_{j}+h(Y_{j})-X_{j}|^{2}\right]$
where $\varsigma\in S_{d}$ is a permutation. More generally, for any random variables $X$ and $Y$,
@@
where the inequality follows by considering better couplings beyond the naive coupling $(Y+h(Y),X)$. Allowing for permutations, or more generally, couplings, offers leeway for higher denoising accuracy, but, perhaps more importantly, it also opens up new theoretical and practical perspectives. In theory, it uncovers new mathematical structures from advanced combinatorics [Com74], optimal transport [Vil08], and higher-order score functions [Bob24]. In practice, for several modern applications in generative modeling, such as image denoising via score-based diffusion models [Son+20, Soh+15, LDK26], the Wasserstein metric is more relevant than the MSE; practitioners care about the quality of the denoiser at the distributional level, not just at the data level.
Traditional data-level shrinkage methods including Bayes-optimal and empirical Bayes denoisers [EM73, JS61] often over-shrink the distribution. They produce denoised distributions that are overly concentrated and do not match the actual signal distribution [GS24, JIS25, Lia25]. We refer the reader to [Lia25, JIS25] for examples and discussions of the benefits of distribution-level shrinkage compared to traditional data-level shrinkage. This paper continues along the path of [GS24, JIS25, Lia25] in studying classic signal denoising through the lens of optimal transport.
@@ 1.2 A Hierarchy of Optimal Transport Denoisers
In this paper, we study signal denoising under the Wasserstein metric and examine whether shrinkage aids denoising in a distributional sense. Recall that for $r\geq 1$, the $r$-Wasserstein metric [Vil03, Theorem 2.18] on the real line takes the closed-form expression,
$W_{r}^{r}(P,Q)=\int_{0}^{1}\left|F^{-1}(t)-G^{-1}(t)\right|^{r}\,\mathrm{d}t$
@@
Considering the denoiser of the form $Y+h(Y)$, we look for denoising function $h:\mathbb{R}\rightarrow\mathbb{R}$ that achieves significantly smaller error than the trivial denoiser $Y$, measured by the Wasserstein metric, namely,
@@
This is analogous to Stein’s paradox, but now at the distributional level. Under mild regularity conditions, the optimal denoiser in the Wasserstein sense is the unique optimal transport map pushing $Q$ onto $P$ [Vil03], denoted by $T_{\infty}:\mathbb{R}\to\mathbb{R}$:
@@
To see how such a hierarchy is constructed, we note that for a monotonically increasing map $y\mapsto y+h(y)$,
$W_{r}^{r}\big(\mathcal{L}(Y+h(Y)),\mathcal{L}(X)\big)$ $=\int_{0}^{1}\left|(id+h)\circ G^{-1}(t)-F^{-1}(t)\right|^{r}\,\mathrm{d}t$
$=\int_{\mathbb{R}}\left|y+h(y)-F^{-1}\circ G(y)\right|^{r}q(y)\,\mathrm{d}y\;.$
@@
The structure of these formulas is governed by a combinatorial relationship involving Bell polynomials; see Theorem 3.2 for the general formula for $h_{k},k\geq 1$:
$h_{k}=$ $-(-1)^{k}\frac{G^{(2k)}}{G^{(1)}}-\sum_{i=1}^{k-1}(-1)^{k-i}\binom{k}{i}\sum_{j=0}^{i-1}\frac{G^{(2k-i-j)}}{G^{(1)}}B_{i,i-j}(h_{1},\ldots,h_{j+1})$
$\hskip 18.49988pt\hskip 18.49988pt-\sum_{j=0}^{k-2}\frac{G^{(k-j)}}{G^{(1)}}B_{k,k-j}(h_{1},\ldots,h_{j+1})\;.$
Here $B_{n,k}$’s are the partial Bell polynomials [Com74] as in Definition 2.3, capturing the combinatorial structure of integer partitions. The above recursions can be solved sequentially to obtain $h_{k},k\geq 1$.
We emphasize that such denoising functions $h_{k},k\geq 1$ only depend on higher-order score functions of $Q$, and are agnostic to the specific form or knowledge of the signal distribution $P$. This phenomenon motivates the name agnostic denoisers. This is in the spirit of the James-Stein denoiser, which is also agnostic to the prior distribution $P$; however, the James-Stein denoiser is only agnostic for discrete distributions with $d$ atoms. In contrast, our denoisers $T_{K},K\geq 1$ are agnostic across all distributions on $\mathbb{R}$.
@@
The paper is organized as follows. As a warm-up for the technical combinatorial structure, Section 2.1 and Theorem 2.4 study the infinite expansion of the optimal transport map $T_{\infty}=F^{-1}\circ G$ in terms of higher-order score functions of the signal distribution $P$. Section 2.2 and Theorem 2.6 quantify $W_{r}(T_{K}\sharp Q,P)$, the denoising accuracy of $T_{K},K\geq 1$, as well as the uniform error of the map $T_{K}$ in approximating the optimal transport map $T_{\infty}=F^{-1}\circ G$. Section 2 is reminiscent of g-modeling in empirical Bayes [Efr14], where one uses prior information to construct denoisers. Though the results in Section 2 are primarily of theoretical interest, they prepare the technical tools for the main results in Section 3.
Moving on to the main results, Section 3 derives and studies optimal transport denoisers via higher-order score functions of $Q$, directly identifiable and estimable based on observations $Y$. Section 3.1 and Theorem 3.2 represent $\{h_{k}\}_{k=1}^{\infty}$ as polynomials of higher-order score functions $\{\tfrac{q^{(m)}}{q}\}_{m=1}^{\infty}$. This result reveals how higher-order score functions encode information of the optimal transport denoising map, a finding new to the literature. We then study estimation strategies and rates of convergence for these higher-order scores, based on $n$ i.i.d. samples $\{Y_{i}\}_{i=1}^{n}$. We present two approaches with theoretical guarantees on the estimation accuracy: (i) a plug-in estimation approach via Gaussian kernel smoothing that estimates $q^{(m)}(y)$ locally at any given $y$, see Section 3.2 and Theorem 3.5 for the rate of convergence; (ii) a direct estimation approach via higher-order score matching that estimates the function $\tfrac{q^{(m)}}{q}$ globally, see Section 3.3 and Theorem 3.7 for the rate of convergence. This is reminiscent of f-modeling in empirical Bayes [Efr14].
@@ 1.3 Relations to the Literature
The optimal transport (OT) denoisers are fundamentally different from the Bayes-optimal or empirical Bayes denoisers [EM73, Rob64, JS61] that minimize the mean squared error. Indeed, the Bayes-optimal denoiser takes the form
@@
different from the hierarchy of OT denoisers $T_{K},K\geq 0$: with $T_{0}(y)=y$, and
@@
Denoiser $T_{1}$ has been studied extensively in diffusion models; see [LDK26, BB25]. [Lia25] studies these two OT denoisers $T_{1},T_{2}$ in a general $d$-dimensional setting, and establishes that these two denoisers are universal denoisers. They remain universal to a wide range of noise distributions, extending beyond the Gaussian setting that requires Tweedie’s formula. On the one hand, this paper generalizes the results in [Lia25] and derives a complete hierarchy of OT denoisers for arbitrary order $K$. On the other hand, this paper considers only $d=1$ and the Gaussian noise.
@@
Empirical Bayes methods [EM73, Rob64, JS61, Gho+25] study the Bayes-optimal denoiser constructed based on i.i.d. samples $\{Y_{i}\}_{i=1}^{n}$ drawn from $Q$. There are two distinct approaches, called f-modeling (modeling on the observed data, the $Y$ space) and g-modeling (modeling on the unobserved prior/signal, or the $X$ space) [Efr14]; we refer the reader to [IS25, Chapter 6] for detailed discussions. [Efr14] states that g-modeling on the $X$ space has dominated the theoretical literature in empirical Bayes, see [Efr16, JZ09, Lai78] for examples. In particular, [JZ09] studied nonparametric maximum likelihood estimation (NPMLE) of the prior distribution $P$ based on i.i.d. samples $\{Y_{i}\}_{i=1}^{n}$ drawn from $Q$. They established the rate of convergence of NPMLE $\widehat{P}$ to the true prior $P$ in terms of the Hellinger distance between the induced densities on the $Y$ space. [Gho+25] draws the connection between score matching [HD05] and SURE [Ste81, XKB12] to propose a SURE training procedure; they also establish the corresponding rate of convergence to the true prior measured by Fisher divergence.
We note that the theoretical empirical Bayes literature primarily focuses on estimating the prior $P$ first, then plugging in Tweedie’s formula to construct the empirical Bayes denoiser. However, as noted in [Efr14], f-modeling directly on the $Y$ space is more natural for constructing denoisers, since the data $\{Y_{i}\}_{i=1}^{n}$ are directly observed. [Efr14] proposed several parametric models on the $Y$ space to estimate the Bayes-optimal denoiser.
@@
This paper takes a fundamentally different approach from the empirical Bayes literature: we construct denoisers directly on the observational $Y$ space via higher-order score functions, without estimating the prior $P$ on the $X$ space. The denoisers $T:\mathbb{R}\rightarrow\mathbb{R}$ constructed in this paper are optimal in the Wasserstein sense, and when applied to $Y\sim Q$, the denoised distribution $T\sharp Q$ matches the signal distribution $P$ up to high accuracy. In other words, this paper proposes directly estimating the optimal transport map $T$ that pushes $Q$ onto $P$, without the knowledge or estimation of $P$. This is in sharp contrast to the g-modeling approach, which first estimates $P$, then uses that information to construct empirical Bayes denoisers. Our denoisers are nonparametric in nature, agnostic to the specific form of the prior distribution $P$. In terms of estimation of the higher-order score functions, we present two approaches: (i) a plug-in estimation via Gaussian kernel smoothing; (ii) a direct estimation approach via higher-order score matching. Both approaches come with theoretical guarantees on the rates of convergence. Crucially, our optimal transport approach discovers and isolates a notion of agnostic denoisers: the denoisers $T_{K},K\geq 1$ are agnostic to the specific form of $P$; they only depend on the higher-order score functions of $Q$.
@@
Score matching techniques have received considerable attention in the machine learning literature [HD05, Hyv07, Vin11, SH19]. The traditional score matching focuses on estimating the first-order score function $\tfrac{q^{(1)}}{q}$ [Hyv07, Sar+18, Sri+17, Fen+24] via Fisher divergence. [Bob24] recently studied Fisher-type information induced by higher-order score functions $\tfrac{q^{(m)}}{q},m\geq 1$. We establish estimation strategies and rates of convergence for these higher-order scores, and derive through advanced combinatorics that these higher-order scores encode full information to construct the OT denoising map with $T\sharp Q=P$.
To our knowledge, this paper is the first to fully characterize the optimal transport denoisers in terms of higher-order score functions; we make precise the connection among Fisher-type information (higher-order scores), optimal transport (Brenier map), and combinatorics (integer partitions) in this classic signal denoising problem.